Mathematics 122 – Gateway Exam – Fall 96

The Gateway exam will consist of ten randomly chosen problems about integration and differentiation. To pass the examination you must do at least nine problems without any mistakes. You need not simplify your answers, but if you do, the simplification must be error-free. If you fail to pass the exam you will be allowed to retake it. When you successfully complete the gateway exam, you will receive 100 points toward your final grade. If you are unable to pass the examination you will not receive any points.

Rules

1. The gateway exam will not be a timed test; you may take as long as you need.

2. You can only take the exam once per day.

3. You will not be allowed to use notes, books, or a calculator.

4. The last day to take the test is October 18

Where and when to take the test

Location: 324 Strong

Times: Monday through Thursday, 7:30am – 7:30pm Friday, 7:30am – 4pm Saturday, 9am – 10:30am

Requirements: You must show you KU ID and know the name of your Instructor (Rodolfo Torres). Ask for 122 Gateway.

Important You will not be able to take the exam on these days September: 9,10, 11, 12, 25, 26, 30. October: 1,9,10,14, 15.